UT Feb 07 2022 - Split Night PETS/DEC



Daytime

  • Meeting at noon - Zoom2 to discuss powering up other instruments for tonight and Monday

Staff

  • OSA: D.G. Huerta
  • ISp: Mtn
  • Mtn Manager: M.Wagner
  • ISA: B. Rothberg, assisted by R.T. Gatto
  • AO: G. Taylor

Night Info

  • Sunset 17:57
  • Nautical Twilight Ends 18:52
  • Astro Twilight Ends 19:21
  • Reconfigure to PEPSI for PETS-4 23:25
  • Moonset 23:53
  • PETS-4 23:55-04:16
  • Reconfigure from PETS to LUCI or MODS 04:17
  • Astro Twilight Starts 05:46
  • Nautical Twilight Starts 06:15
  • Sunrise 07:10

Night Plan

If LBTI is not online or LBTI conditions not met

  • Configure for LBCs at handover
  • LBC/TMS Testing twilight-23:25
    • Any conditions
    • TMS Setup
      • Service_Observing/2021B_D_E/TMS_13Dec
      • ssh telescope@robs
      • cd /lbt/lbto/supportscripts/MetrologySupport
      • Check if reference file is copied over (J. Hill or O. Kuhn will verify)
      • start ./tms_loop.pl to get mirrors into position
        • When running, bring up the PSF GUI, Primary and Position Details - look at Group Rotation - shifts from away 0 then we can stop and move to co-point field for dofpia
        • This may take a while
      • CTRL^C the TMS
      • Go to co-point field, run dofpia (not dohybrid)
      • AFTER dofpia is completed then run TMS
        • First Set reference: ./lbc_set_tms_offset.py -s 2
          • this will take a while before returning prompt
        • In MetrologySupport directory execute ./tms_loop.pl
          • if one passive and one active - then ./tms_loop_right.pl for DX(Red) or ./tms_loop_left.pl for SX(Blue)
      • co-point with lbcrangebal, then go to field of choice
      • Start running scripts
      • CTRL-C before running DOFPIA again - if run DOFPIA and TMS simultaneously, the two will "fight" - TMS will backout anything DOFPIA does!
    • Scripts; ~/Service_Observing/2022A_D_E/LBC/TMS_Active9/TMS_ActPass9_OT.xml
      • field_0400+25_30min.xml --- science script

      • field_0400+25_CP.xml --- copointing script

      • field_04+25_FC.png --- copointing finding chart

      • Note that two fields should be used tonight
      • For the beginning of the night: 06 30 00.0 +23 00 00 (ecliptic field) 1mn exposure time (g-r as usual)
      • Beyond 21:00 (local time): 06 27 53 +22 30 50 (C/Halley) 2mn exposure time (g-r as usual) - note that we stay in sidereal motion to use images for IQ - Halley is TNO-like, moving very slowly

If LBTI is available

  • Configure for LBTI at handover
  • LBTI - CQ Tau w/ALES+vAPP coronagraph sunset-23:25
    • Clear (photometric)
    • Seeing < 1.2"

If LUCI-AO not available

  • Reconfigure from PETS/PEPSI to MODS 04:17
  • MODS-science 04:18-twilight

If LUCI-AO (or LUCIs) available

  • Reconfigure from PETS/PEPSI to bino-LUCIs (or LUCI-1 if LUCI-2 not powered on) 04:17
  • LUCI-1 AO Testing
    • /home/lbto/Service_Observing/2022A_D_E/LUCI/AO_ESM/
    • ESM OT: ESM_V32.xml
    • AO OT: AO_Testing_V58.xml
      • OSCO
      • Bin 1 Full - highest priority (OSURC science will need)
      • Bin 3 Full
      • Bin 1 "cool" target
      • ESM Bin 1
      • ESM Bin 2
  • If AO not available or poor weather conditions
    • Local Mergers - bino-(or mono) LUCI - Rothberg
      • /home/lbto/Service_Observing/2022A_D_E/PoorWeather/Rothberg/LocalMergers/COBandhead_Mergers_V37.xml
      • Any conditions
        • UGC 9829
        • UGC 10675
        • UGC 10607
        • NGC 4004

Night Log

Closed Dome

  • UT 00:14 - bringing up LBC GUIs and powering up
  • UT 00:18 - O.Kuhn has already coped over the refernce file for TMS for tonight - 20220127 into /tmp
  • UT 00:32 - powered up - taking test biases
  • UT 00:23 - bringing up LUCI GUIs - running as observer on OBS5 to test RTDs
  • UT 00:25 - running init all on LUCIs - taking 2 test darks 2.51sec
  • UT 00:26 - LBC biases look OK
  • UT 00:27 - test darks LUCIs look OK - putting N30Fieldstop in for LUCI-1
  • UT 00:30 - checking N30 Fieldstop alignment
  • UT 00:31 - duh might be more useful if the N30 camera is used too....
  • UT 00:34 - L1 0003-0006 - completed alignment putting blinds back in - LUCIs ready for AO tonight
  • UT 01:09 - opening dome for LBTI

LBTI

  • D-time target
  • Seeing 1.6" in setup field....eeek!
  • Testing safe skip and replicating problems
  • UT 02:15 - stopping and handing over to TMS/LBCs

LBC

  • UT 02:17 - H.Choi powering up lasers
  • UT 02:20 - lasers powered up and ready
  • UT 02:34 - connecting LBCs to TCS
  • UT 02:38 - runing tms_loop.pl - to get them into position (running both)
    • When running, bring up the PSF GUI, Primary and Position Details - look at Group Rotation - shifts from away 0 then we can stop and move to co-point field for dofpia
  • UT 02:41 - TMS shows pose calculation failed in Terminal - something wrong?
    • * pose calculation failed continuously (15) times, aborting... *
  • UT 02:44 - stopped the loop - H.Choi is rebooting the PC - will need about 5minutes
    • If the channels are different between reference and whats working tonight - will need to run DOFPIA - and then set a new reference
  • UT 02:52 - starting script again for TMS - no error messages and Group rotation changes
  • UT 02:54 - sending co-point script - track preset recieved - pressed stop - will run dofpia NOT dohybrid
  • UT 02:56 - reaching limits on X and Y on M1 SX - close on DX M1
  • UT 02:58 - arrived at field - going to restart PCS to see if this clears errors
  • UT 03:02 - PCS restarted, initialized M1 SX
  • UT 03:04 - moving off target to zenith - initialized both primaries
  • UT 03:05 - starting active tms script
Mystery Field
  • UT 03:08 - stopped script - group rotations updated -moving to MysteryField co-point
  • UT 03:10 - running dofpia
  • UT 03:13 - seeing on DIMM~2.5"
  • UT 03:18 - DOFPIA finished - lcollimation ooks worse
  • lbcrangebal_mysterfield_copoint.png
  • ~1 degree differnce face-to-back temperature difference
  • UT 03:24 - Runing DOFPIA again to set a new reference
  • DOFPIA converged after 1 iteration
  • UT 03:30 - running DOFPIA again - DIMM 3.3"
  • UT 03:36 - added +1500 Z7 - re-running DOFPIA - see if it kicks it in the right direction
  • UT 03:38 - DOFPIA finished up - removed Z7
  • UT 03:41 - ./lbc_set_tms_offset.py -s 2 - set reference point
  • UT 03:44 - Now run script ./preppupils.py - more interested in thermals and relative changes
    • should see Z4 offset in PSFGUI Z4=-1050 global and Z4=-875 in global
    • running ./tms_loop.pl
    • Running MysteryField script
    • Changed mystery field script to 32sec exposures x 200 exposures g-sloan and r-sloan
  • UT 03:50 -started updated script
    • will transit at 80.3 deg at 21:39 local or 04:39 UT
  • Images actually look in-focus - as opposed to extra-focal, instrument offset is now -0.6
  • UT 03:54 - pausing - change from -0.38 Z global offset to -1.18 mm in Z(mm) position - new total is -1.251 for SX and -1.447 for DX
  • UT 03:57 - this seemed to be undone by something - TMS?
  • UT 04:04 - created a version of RB_fastextra - RB_fastextra4.ob with N=200 and 32sec in current directory
  • UT 04:09 - manually added z7 +1500 - coeffiecient is now -1123
  • UT 04:11 total RX collimation is -35.7 on DX, now added +3000 z7 global offset, RX collimation changed to -49.499 - total z7 coefficient DX = 376
  • UT 04:13 - RX collimation now -49.57 - TMS didn't back out the big change
  • UT 04:16 - add +500 z11 on DX - total z11 now 1092.2
  • UT 04:18 - add +500 z11 on DX - total z11 now 1592.2
  • UT 04:19 - add +200 z11 on DX - total z11 now 1792.2
  • UT 04:21 - add +200 z11 on DX - total z11 now 1992.2
  • UT 04:22 - add +300 z11 on DX total z11 now 2292.2
  • Plan is to just collect the extra-focal images until handover given the terrible seeing
  • UT 04:40 - RX collimation hasn't changed by much - now -50.52
  • UT 05:02 - DX face-to-back temp changed - noticed more coma on the LBC-Red images now
  • UT 05:18 - changed z4 by -10000
  • UT 05:19 - changed by z4 by -3000, new z4 is -13175
  • UT 05:22 - changing z4 on SX to -6550 global offset
  • UT 05:24 - central hole on blue is no longer visible
  • DIMM says up to 4" seeing
  • UT 05:49 - DIMM shows 5"!!!
  • UT 06:01 - DIMM has calmed down to 4.1"... winds are picking up - some gusts to 16-18 m/s - ambient temp started at -5C now -8.3C
  • UT 06:07 - humidity has risen as well, now 65% range - started night much lower
  • UT 06:23 - LBC-R images can barely see the central hole
  • UT 06:25 - stopping LBCs and turning of TMs
  • UT 06:26 - powering off LBCs

PEPSI/PETS-4

  • UT 06:25 - reconfigure for binocular PEPSI
  • Target is TOI 561
  • Having trouble acquiring - seeing so poort
  • Closing dome - gust hit 21.3 m/s
  • Attempt to collimate elsewhere before closing? - if DGHuerta can collimate away from gusts, then go back to field will stay open
  • UT 07:27 - trying a brighter star to collimate on
  • Suggested turning off zernikes and just putting star in pnhole, seeing is > SH spots - trying to collimate is meanngless
  • UT 07:55 - DGHuerta calling J.Hill
  • UT 08:08 - finally preset to target - in pinhole - starting integration
  • UT 08:32 - fire alarm at summit - lost prest on DX - not sure what's going on...
  • UT 08:36 - I see lights in the enclosure (flashlight) - mtn staff looking for something...
  • UT 08:42 - Air pressure on elevation and azimuth brakes failed - no smoke smelled - no smoke seen
  • UT 08:43 - closing dome for safety
  • UT 08:47 - still investigating
  • RM 408 - location of alarm - no indication of source
  • UT 09:05 - will start up mirror ventilation, air pressure, recover panics on M1s
  • UT 09:07 - smoke sensor from lieber UPS - UPS looks fine
  • Apparently fire alarm didn't trigger call from local fire department - MGIO repeater did send out alert about fire at LBT
  • UT 09:24 - opening enclosure
  • UT 09:34 - left primary panicked again
  • UT 09:55 - opening dome
  • UT 10:02 - sending preset
  • UT10:15 - primary issues again
  • UT 10:32 - on target and taking data
  • UT 10:46 - SX M1 having an issue again - had to call up to let DGHuerta know
  • UT 10:57 - seems to be ok now
  • UT 11:14 - star vanished on SX
  • PEPSI sees 6.67" and 7.37" seeing on AGWS
  • UT 11:16 - PEPSI done reconfiguring for bino LUCIs

LUCI

  • UT 11:16 - DX M1 panicked
OSCO
  • Target will be HD 118701 - RA 13hr Dec +33 (R=7.48) - will point and collimate near star and see if seeing is crud-tastic still
  • G.Taylor is cooling OCAM just for completeness....
  • UT 11:33 - DGHuerta still working on DX M1
  • UT 11:37 - acquire preset for pointing
  • UT 11:41 - active preset for collimation....over or under 4"?
    • FWHM in guider bouncing from 3.2" to 5.2" to 4"
  • UT 11:45 - trying to collimate on a 4th mag star (preset failed on 6th mag)
    • GS lost after 6 cycles according to IIF
  • UT 11:47 - trying R=7.3mag star for GCS....
    • boucing between 2.8"-3.4" on guider - SH spots on WFS look fully illuminated
  • Abandoning AO, will try something SL, G.Taylor warming up OCAM
SL Backup
  • UT 11:51 - authorizing bino-LUCI, will point and collimate....
  • putting N30fieldstop away to save time
  • UT 11:56 - authorize for DX LUCI-2 seems to be failing
  • UT 11:59 - restarting GCS Right
  • UT 12:03 - restarting GCC right again
  • UT 12:05 - restarting GCS right again
  • at the very least we'll point and collimate and make sure things are working since powering up the AGWs
  • UT 12:10 - AGW2 failed with 4"+ on the guider (its working it seems - but it couldn't keep a lock on the star) - star is R=7.3mag
  • I'm calling it ... brightest GS I had was R=9.... plus...we never actually collimated so who knows how out of focus the image would have been on the focal plane...
  • UT 12:20 - closing dome and DX panicked

-- BarryRothberg - 05 Feb 2022
Topic revision: r4 - 07 Feb 2022, BarryRothberg
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback